# Inference-time Alignment in Continuous Space

### v4

@@ 5.6 Further Analysis
Multi-Dimensional Alignment. Another advantage of our SEA based on EBM is its ability to directly combine probability distributions, enabling compositions of rewards through sampling from the combined energy du2020compositional . In Table 8 in Appendix C.1, we evaluate the capacity of SEA to generalize to reward combinations, demonstrating effectiveness of SEA in achieving multi-dimensional alignment.
@@ C.4 Additional Models, Datasets and Baselines
Method
AdvBench
HarmRate $\downarrow$ (%)
TruthfulQA
TruthRate (%)
TruthfulQA
InfoRate (%)
$H$-RLHF
WinRate (%)
@@
Method
AdvBench
HarmRate $\downarrow$ (%)
TruthfulQA
TruthRate (%)
TruthfulQA
InfoRate (%)
GSM8K
Acc (%)

### v5

@@ 5.6 Further Analysis
Multi-Dimensional Alignment. Another advantage of our SEA based on EBM is its ability to directly combine probability distributions, enabling compositions of rewards through sampling from the combined energy Du et al. (2020). In Table 8 in Appendix C.1, we evaluate the capacity of SEA to generalize to reward combinations, demonstrating effectiveness of SEA in achieving multi-dimensional alignment.
@@ C.4 Additional Models, Datasets and Baselines
Method AdvBench HarmRate $\downarrow$ (%) TruthfulQA TruthRate (%) TruthfulQA InfoRate (%) HH-RLHF WinRate (%)
@@
Method AdvBench HarmRate $\downarrow$ (%) TruthfulQA TruthRate (%) TruthfulQA InfoRate (%) GSM8K Acc (%)